3. A pair of fair dice are tossed. Find the probability of getting (a) a total of 4. (b) at most a total of 4
Answer #1

PS: If the image is not visible, right click on image and select open image in new tab option

Sample Space

Total 36 Outcomes

a)

Probability = Favourable Outcomes/Total Outcomes

Dice 1

Dice 2

Sum

1

3

4

2

2

4

3

1

4

Hence, P = 3/36 = 0.0833 or 8.33%

b)

Dice 1

Dice 2

Sum

1

1

2

1

2

3

1

3

4

2

1

3

2

2

4

3

1

4

Hence, P = 6/36 = 0.1667 or 16.67%